Hours 1001

From DIDEAS Wiki
Jump to: navigation, search

Jan 2010


Mon Jan 4

  • 9:30a - 3:10p -30
    • prepare PSER1eb and PSER2hb
    • layout inverter oscillator and complete two eddy current sensors PCBs
    • panel and submit PCBs for fabrication
  • 4pm - 7:30p -45
    • discuss w/DG LTF status
    • resubmit PCBs after project rejected.
    • biquad filter research and implementation
    • discuss w/RC LTF status (41 hz) and V2 status


  • LTF data analysis
  • get V2 going

tue jan 5

  • 10:30a-3pm -1:40
    • biquad filter tested on linux, added support commands under tkcmdset
    • converted LTF data from Dec30
  • 3:30p - 9:40pm
    • work with GG and CA on IMU programing, hardware diagnosis
    • review LTF data with DG, experiment with LTF

wed jan 6

  • 8am - 11:40a +60min
    • review robo_config for ability to load/save tables
    • develop test procedure for ED / GG to load / save EEPROM tables on Spld.
    • discuss w/DG how to proceed w/LTF experiments and XP software requirements.
    • complete integration of biquad.c into MC firmware, add prototypes to result in cleaner MC compile
    • located tool to design biquad filter coeffs, designed 200 & 300 hz Bstop filters.
  • 3pm - 11:25pm -30 +30
    • run LTF w/DG
    • tested biquad filter
    • update LTF SC/MC
    • ran sine sweeps and white noise.
    • converted data emailed to DG

thur jan 7

  • 15+20+20+1hr
    • discuses w/RC robot problem and patent
    • phone w/DG re LTF results
    • phone w/GG re LTF an robot results
    • read patent proposal

fri jan 8

  • 1 hr
    • working with DG on LTF

sat jan 9

sun jan 10

  • 2pm - 8:45p -30
    • meet w/RC @CKF - discuss robot testing, holster, block diag, belt detector
    • review SERDES
    • discuss SERDES ideas w/MM - decide on FPGA w/1 TX, RX, CLK
    • review of patent
    • consider alternatives to connecting to SC/MC. Possibly HS USB?

mon jan11

  • 4-9pm
    • discuss robot arm control w/ZH and GG w/python
    • discuss w/GG java UDP for plot_lifefix
    • work w/DG on LTF. Run white noise tess, followed by a large series of sine sweeps.

tue jan 12

  • 1.5 hrs - reading patent
  • 3-8:30p
    • sign patent papers
    • review matlab UDP java tool
    • wire up new LTF power supply.
    • work with DG on LTF freq response, and biquad filter.

wed jan 13

  • 1 hr
    • investigate methods of multiplexing ISP for debug deep in the robot.
    • 3 methods:
      • develop our own programmer that MPLAB can use (or copy one)
      • implment the ISP statemachine in the FPGA so it can determiante directionalty
      • use an analog swtich to multiple the PGD signal - so we don't have to understand the protocol, use SER-DES for the remaining signals.
  • 4pm-5:30p
    • cut pcb pannel
    • assemble eddy current sensor PCB and test.
    • ???
  • 6:15pm - 10:15pm
    • work with DG on LTF.
      • verify freq rsp of pole zeros with fvtool
      • calculate new pole zeros
      • test biquad on LTF
      • start run_lifefix_v2, issues with KCA.

thur jan 14

  • 2:35 - 10:20p
    • digikey order for RJ45 magnetics and other items
    • assist ZH with real-time work loop plotting
    • bring up LTF, try to understand commutation problem
    • LTF PID tuning
    • LTF work loop studies.

fri jan 15

  • 2.5 hrs
    • complete excell entry for Dec and to present time for Jan
    • email summary of the cause/solution to the matlab plotting problem
    • email summary of LTF ground loop problems - and a suggested change.
    • documentation for Pser2hb, pser1eb
    • assist w/running of LTF - find the motor isn't moving, likly over heated.
  • 3:30p-11:20p
    • assist CA w/ankle rebuild and programming of SWIFI, Wiport, SC, and MC
    • restart LTF
    • investigate the error in the resistance measurement, find that qId depends in non-linear way on qVd

sat jan 16

  • 10a-12p
    • review LTF, code real-time temperature plotting
  • 1:10p-9:40p -15
    • review GG code merge, program on V2.
    • improvements to matlab plotter, reads a user config file for plot titles / scale, experment with alternative UDP libraries
    • testing MC on rev300 hardware.
    • re-write and test acoustic mods to statecontroller

sun jan 17

  • 3:15p - 11pm
    • investigate the cause of the LTF stopping, appears to be Pser2h non-isolation, fixed.
    • studied the effect of motor rotation on apparent resistance (added tested to benchtest.py)
    • looked for relationship between ADC sampling position and measured current (non found)
    • compared DC type resistance measurement with AC measurement method. Found good agreement.
    • fixed bug in RLFv2 where text output was incorrectly scaled.
    • (start) modify RLFv2 to output data computer readable format.

mon jan 18

  • monitor LTF overnight
  • 2 hrs
    • discuss status w/DG on LTF status, count number of cycles
    • improve RLFv2 to encode exit error for retry*
  • 4pm-11:30p
    • assist GG w/statemachine bug (400mS in PP)
    • continue testing development of RTLFv2
    • successful test LTF w/48V supply
    • w ED/investigate PSER2h baudrate problem, make kernel modules changes. After, there are still an issue setting the baudrate - however multiple settings seem to take effect
    • investigate methods of testing motor automatically with benchtest
    • investigate possibility of using the JR3 for motor force control
    • start process of converting SWIFI data from LTF weekend run

tue jan 19

  • 2 hrs
    • create LTF parts list [1]
    • convert SWIFI LTF weekend data.


  • 24V supply for battery charger
  • 2:30p - 8pm
    • test auto retry code
    • w/ED discuss future calibration and LTF logging
    • get wifi_fast.py running on LTF and sending to matlab
    • 6pm start LTF and monitor progress w/DG
    • LTF stops occasionally due to excessive belt range. has no obvious cause.
    • improvements to retry code to auto-restart
  • 8:50p-11:40
    • start to code the mc_resistance as a stand alone class.
    • improvements to motor temp algorithm and cycle type measurement
    • improvements to retry code to auto-restart
    • the auto-retry will break the total-run time duration

wed jan 20

  • 60 mins :
    • study method of controlling multiple LTF
    • email DG on same
    • discuss K w/JW
  • 3:15-7:15p
    • discuss LTF vibrations and future w/DG
    • meet with HH and JW re new project,
    • discuss SM anomalys w/HH - develop new ES transition
    • discuss SM, PP, etc w/ZH
  • 10:25p-2:20a
    • coded better transition for ES2LS
    • developed method of tracking state transitions, started coding in SM
    • developed virtual spring module
    • wrote program to parse LTF log for import to ML
    • tested python's ability to read JR3 at 1khz, and send a byte over serial port.

thur jan 21

  • 1.5 hrs
    • modify virtual spring code to make use of AA and AAdot, clean up helpper functions
  • 4pm - 1:40a
    • arrive to find LTF has stopped and has reached a hard stop, decided that cause was related to miss commutation or PI sign problem
    • needed to figure out how to decode a very large SWIFI file.
    • added to RLFv2 and -N option to not do commutation
    • added to run_retry the addition of the -N suffix after the first retry
    • review LTF status w/DG
    • discuss TS and exp parameter decay w/RC
    • create an automatic new file name class for wifi_fast, every 10 mins a new log file is used.

fri jan 22

  • 1 hr
    • admin
    • email documenting LTF logs, locations, etc
    • email questioning hall sensor and calibration procedure
  • 5:15pm - 7:40pm
    • w/GG on robo config, explain MILE related MC changes since Oct 13 split
    • modifications to do_command.py to support system wide install
    • discuss calibration requirements w/GG and ED - locate matlab script to create python calibration table.
    • discuss repair of S2 and options, decide best to order more Rev200 boards, create build ECOs same
    • tested new 24V supply

sat jan 23

  • 4:30p - 9pm -1.5
    • review LTF data from jan15-17 session
    • modify wifi_fast smart log file class to optionally use threading
    • install SWIFI on S3.
    • began experimentation with latest SC/MC on S3/V2


  • determine why pser2h and swifi FUP communications collide - even with USB off.

sun jan 24

  • 1 hr
    • assembly docs for LTF
    • admins
  • 2:15p-8:45p
    • test Rev300 PCB w/Rev200 SC/MC on bench
    • wire 48V LTF power supply
    • review WIP4c and PSER2h schematics - collision is the result of both devices using 1k protection resistors.
    • test latest firmware, correct bugs.
    • correct new log method problems w/wifi_fast. adds logs to zip file periodically - works well.


  • test rev 300 PCB on LTF
  • get robot working
  • test VS code
  • develop current offset calibration

mon jan 25

  • 1.25 hr - email, discussion re LTF power supply, and admin, email to CA re need experiment with PSER mods.
  • 4:05p - 12:45a (-1 hr)
    • assist CA w/EC30 resistance measurement, motor actually works
    • sketch schematic for LTF power supplies
    • meet with Charley doing simulink MC and BC models
    • testing S3 - debugging code
    • test V2, seems to lack working encoder


  • modify / test Pser2h with smaller Tx protection resistors.

tue jan 26

  • 2 hrs
    • email re MPD build of new R200 boards, need to resolve issue with PSER1b
    • submit NOV 09 invoice
    • work on BOM for PSER1eb, PSER2hb.
  • 2:30-7:15p
    • discuss w/ED force ripples while calibrating some robots.
    • decided to make a true open leads mode in the MC firmware (including 751x) - added MODE10 which does exactly that. This is in addition to the LTFs open leads in MODE0.

wed jan 27

  • 10:50-12:15
    • email re modifications to FET boards for safety fets., update WIKI, looking for IW Wiki, admin
  • 1p-7:15p
    • restart LTF
    • meet with JW, HH, RC, DG, CW re knee design
    • assist CA with testing of Rev200 boards.
    • disuss w/GG getting MPD moving on Rev200 boards - send an email


  • remove safety fet resistor from REV300

thur jan 28

  • 11:30a - 12:35p
  • 4:15pm -12a
    • fix read of encoder bug in MC that could result in 1 rev error
    • modify RLFv2 to require 2 successive belt range errors before exit
    • assisted ED and ZH w/cause of series spring calibration anomaly.
    • discuss creation of new PVT and thrust vectors w/DG and ZH.
    • discuss w/GG addition of hall table to robo_config.
    • work on V2 and firmware. Discover SG wiring problem.

fri jan 29

  • 30 mins home
  • 1:00p - 1:45p
    • get an update on the pyramid
    • verify test procedure to use do command to read pyramid and to plot @500hz
  • 3:30p - 5pm
    • Pser BOM
    • V2 firmware debug
  • 6:30p-9pm
    • meet w/RC and ZH to discuss motor anomaly events, make punch list
    • test V2 pyramid, TQ now working
    • apply MC overload, encoder to S3 firmware
    • investigate windup issue, LTF is not using modified MC filter.

sat jan 30

  • 1 hr
    • review LTF status, create plots analysis
  • 2:15pm - 6pm
    • experiments with motor PI filter and over-windup. Was able to verify the current spike associated with a sudden stop. Make changes to SC/MC firmware, and LTF to always turn on the KC term.
    • should use motor speed to supplement
    • replace 24V LTF supply and test

sun jan 31

  • 10:30 hr - 12:30
    • review LTF status, create plots analysis
    • figure out how to unzip a unreadable 6gb zip file (take the head, and then repair)
    • compare workloops from the start of run and present time.
  • 4:15p-10:15p
    • review LTF, estimate rate of backlash growth.
    • assemble LTF power supply kit.
    • apply and test SC and MC patches to the Spld3 firmware.
    • tested the new V2 pyramid. Has signal on both TQ and LD.